# Towards smart and adaptive agents for active sensing on edge devices

### v1

@@ V-B3 System
Integrating the perception and planning modules produced a highly efficient saccade agent capable of real-time performance. Optimized with TensorRT FP16, the perception module processes frames at a rate of 45 FPS, enabling rapid feature detection. Complementing this, the active inference module operates at 4 FPS for planning and decision-making. The overall system presents, as shown in Table I, a lightweight architecture of just 2.3 million parameters and a minimal memory footprint under 1 GB, including all necessary libraries and runtimes. This configuration balances high-speed perception and adaptive decision-making, allowing the agent to adjust its focus dynamically on constrained resource devices.
Figure 5: Latency of the main functions of the active inference’s module for planning when deployed on the Nvidia Jetson Orin single-core CPU with the PyMDP library and JAX backend.
@@
2.4 0.94 45 4
TABLE I: Saccade agent’s system parameters on the Nvidia Jetson Orin.
@@ V-C Discussion
A key challenge in balancing performance and accuracy on edge devices for active sensing lies in the trade-off between perception and planning modules. While the perception module shall provide low latency to detect the movement of people or objects in the image, demanding high compute and memory requirements, the active inference planning module should conserve computational and memory resources, running at a lower frequency as camera movements are less frequent and slower in IoT surveillance and robotic applications.
Edge devices’ limited memory and computational power highlight the need for innovative approaches to optimizing the planning and perception modules. Developing smaller and more efficient algorithms that work within these constraints is essential for improving overall system performance. Following this principle, we propose an active inference module that, in contrast to massive foundation LLMs that employ billions of parameters to achieve generalization, our model only contains 1.34K parameters, achieving effective planning and adaptation on resource-constrained devices.
@@ VI Conclusions and Future work
This work has introduced a smart edge agent composed of a deep learning-based perception module and an active inference planning module for active sensing. The system demonstrates the feasibility of adaptable, on-device surveillance and robotic solutions and their potential to effectively handle real-world challenges.
Our study highlights the potential of active inference in edge-based systems. Active inference offers a powerful framework that accounts for the environment’s underlying dynamics and inherent uncertainty, allowing the agent to actively reduce ambiguity in its field of view or explore the environment for new information. This approach not only enhances the functionality but also points to new possibilities for adaptive, resource-efficient decision-making in IoT and edge computing applications.
In future work, we aim to optimize the deployment of the active inference module, enhancing its efficiency and scalability to address more complex scenarios. In addition, we plan to compare the proposed system to other state-of-the-art works and benchmark it on popular challenges, such as the Habitat Navigation Challenge. Finally, extending this edge deployment workflow to other active inference-based models could unlock further applications, enabling a new generation of adaptive edge devices capable of context-driven interactions in complex environments.

### v2

@@ V-B3 System
Integrating the perception and planning modules produces a highly efficient saccade agent capable of real-time operation with only 2.3 million parameters. Table 1.1 summarizes key configurations, illustrating the trade-offs between latency and memory consumption. For example, in configuration A, the perception module optimized with TensorRT-FP16 achieves a throughput of 45 FPS, enabling rapid feature extraction. In parallel, the active inference module using TFLite can perform planning and decision-making at up to 250 FPS, with a total memory footprint under 1 GB. The decision-making rate can be reduced to 108 FPS using ORT, lowering memory usage to 533 MB, as shown in configuration B. Finally, configuration C employs the ORT (intra-operator parallelization) for perception, decreasing the perception rate to 9 FPS while providing a compact memory footprint of 304 MB, including all required libraries and runtimes. This configuration effectively balances perception and decision-making’s latency and memory, enabling dynamic adaptation and efficient operation on resource-constrained edge devices.
Figure 7: Planning module memory profile of the active inference model when executed with different deployment inference engines on the Nvidia Jetson Orin NX CPU and GPU.
@@
A   2.3    947   45    250
B   2.3    533   45    108
C   2.3    304    9    108
TABLE I: System parameters: Saccade agent deployed on the Nvidia Jetson Orin NX
@@ V-C Discussion
These results highlight the effectiveness of combining deep learning for perception with active inference for planning in edge-based intelligent agents. With only 1.34K parameters, the active inference module achieves real-time planning while maintaining minimal computational and memory demands. In contrast to large-scale foundation models that rely on billions of parameters to achieve general-purpose intelligence, our domain-oriented approach demonstrates how compact, task-specific active inference models can complement deep learning perception. The adaptability of our edge-deployment workflow illustrates deployment interoperability, facilitating the exploration of diverse configurations to achieve a suitable balance between latency, accuracy, and resource efficiency.
@@ VI Conclusions and Future Work
This work has introduced a smart edge agent composed of a deep learning-based perception module and an active inference planning module for active sensing. The system demonstrates the feasibility of adaptable, on-device surveillance and robotic solutions and their potential to handle real-world challenges effectively. Our study highlights the potential of active inference in edge-based systems. Active inference offers a robust framework that accounts for the environment’s underlying dynamics and inherent uncertainty. It allows the agent to actively reduce ambiguity or explore the environment for new information. This approach establishes a foundation for efficient edge agents for adaptive, resource-efficient decision-making in IoT and edge computing applications.
In future work, we aim to compare the proposed system against other state-of-the-art works, showing a quantitative functional comparison on popular challenges, such as the Habitat Navigation Challenge. Finally, we envision extending this concept to other applications, which could unlock a new generation of adaptive edge devices capable of context-driven interactions in complex environments.